Circular No. 7252

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999do IN MARKARIAN 922 D. C. Leonard, R. T. Chornock, and A. V. Filippenko, University of

California at Berkeley, report that a CCD spectrogram (range 330–1000 nm), obtained on Sept. 10 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 1999do (cf. IAUC 7242) is of type Ia. Prominent absorption lines of Ca II (H and K) and Si II (635.5 nm), corrected by the redshift of the parent galaxy (6400 km/s), imply expansion velocities of ∼ 13 000 and 11 000 km/s, respectively. A type-Ia spectral aging algorithm (Riess et al. 1997, A.J. 114, 722) gives an age of 9 days past maximum.

COMET C/1999 J3 (LINEAR)

Ephemeris from orbital elements on MPC 35553:

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 9 8^{h}34.^{m}85 +46^{o}29.^{′}3 1.207 0.997 52.^{o}6 53.^{o}4 9.4
14 8 29.48 +43 10.7 1.098 0.983 55.5 57.5 9.1
19 8 23.72 +39 05.3 0.985 0.978 58.9 61.6 8.9
24 8 17.50 +33 53.7 0.871 0.980 62.6 65.3 8.6
29 8 10.64 +27 07.8 0.761 0.989 66.7 68.4 8.4
Oct. 4 8 02.82 +18 08.8 0.660 1.007 71.3 70.3 8.1
9 7 53.52 + 6 13.5 0.576 1.030 76.5 70.5 7.9

Visual m_{1} estimates: Aug. 25.08 UT, 9.5 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); Sept. 6.07, 8.9 (J. Kysely, Vlasim, Czech Republic, 7×50 binoculars); 8.08, 9.0 (B. H. Granslo, Fjellhamar, Norway, 0.07-m refractor); 9.10, 9.6 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 10.88, 10.0 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector).

COMET C/1995 O1 (HALE-BOPP)

Ephemeris from orbital elements on MPC 35204:

1999TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 9 7^{h}02.^{m}82 −68^{o}48.^{′}8 9.329 9.300 85.^{o}3 6.^{o}2 12.5
19 7 07.07 −69 56.5 9.415 9.375 84.7 6.1 12.6
29 7 09.7 −71 06.4 9.502 9.449 84.0 6.1 12.6
Oct. 9 7 10.3 −72 16.9 9.589 9.523 83.2 6.0 12.7
19 7 08.6 −73 26.4 9.677 9.597 82.4 5.9 12.8
29 7 04.2 −74 32.8 9.765 9.670 81.6 5.8 12.8
1999 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green